### v5

@@ Appendix D The behaviour of the specific heat
As mentioned in our main text and explained in Appendix C, the specific heat is used to define the cutoff of the spectrum of operator $\mathbf{L}^{\left(q\right)}$, which helps determine relevant scales in the system. Specifically, the $q$-order specific heat in Eq. (23), derived using the first derivative of the $q$-order spectral entropy, is used to define the optimal time scale, $\widehat{\tau}^{\left(q\right)}$, of the SRG in Eq. (24). Then, the inverse of time scale $\widehat{\tau}^{\left(q\right)}$ is utilized to define the number of characteristic modes for describing the system near criticality (see step (2) in the SRG).
Figure 9: The simplex renormalization group (SRG) with different $\left(p,q\right)$. (a) We generate a system with $200$ units whose pairwise interactions follow a Watts-Strogatz network (each unit initially has $5$ neighbors and edges are rewired according to a probability of $0.05$) [80]. Then, the SRG is applied to renormalize the system on the $p$-order according to the properties of $q$-order interactions ($p=1$ and $q=2$). The first five iterations of the renormalization flows on $q$- and $p$-orders are presented, accompanied by their corresponding high-order path Laplacian operators. (b) For comparison, we show the renormalization flow of the system on the $p$-order guided by the $q$-order when $p=q=1$.
@@
Appendix E Comparison of the SRG across different settings of $\left(p,q\right)$
In Fig. 9, we use the SRG to process a system whose pairwise interactions exhibit small-world properties [80].
We first renormalize the system on the $p$-order under the guidance of $q$-order interactions ($p=1$ and $q=2$). As shown in Fig. 9(a), integrating out short-range $q$-order interactions lead to an organizational-structure-preserving renormalization flow of $p$-order interactions. Because high-order interactions are usually distributed in a sparser manner than pairwise ones (e.g., see the flow of $\mathbf{G}^{\left(q\right)}$), the renormalization flow guided by $q$-order interactions reduce the system less drastically.
For comparison, we renormalize the system on the $p$-order according to the properties of $p$-order interactions themselves (i.e., this is equivalent to setting $p=q=1$ in the SRG) in Fig. 9(b). The generated renormalization flow reduces the system more significantly because $p$-order interactions, including those short-range ones, are distributed in a dense manner. The differences between Fig. 9(a) and Fig. 9(b) qualitatively suggest the effects of higher-order interactions on lower-order interactions during renormalization.
{adjustwidth}
-0cm
Figure 10: The existence of the multi-order scale-invariance (Part I). (a-e) The degree distributions and the Laplacian eigenvalue spectra of different synthetic interacting systems under the transformation of the SRG are illustrated after being averaged across all replicas. The multi-order Laplacian operator in Eq. (1) is applied to implement the SRG.
@@ Appendix E Comparison of the SRG across different settings of $\left(p,q\right)$
Appendix F Scaling relation of the Laplacian eigenvalue spectrum in the ergodic case
@@
Appendix G Scaling relation of the Laplacian eigenvalue spectrum in the non-ergodic case
@@
where $\Lambda^{\left(q\right)}_{k}$ and $\Lambda^{\left(q\right)}_{k}\left(i\right)$ denote the eigenvalue sets of $\mathbf{L}^{\left(q\right)}_{k}$ and $\mathbf{L}^{\left(q\right)}_{k}\left(i\right)$, respectively. According to Eqs. (57-58), the characteristic polynomial roots of $\mathbf{L}^{\left(q\right)}_{k}$ can be constructed using the combination of the characteristic polynomial roots of $\{\mathbf{L}^{\left(q\right)}_{k}\left(1\right),\ldots,\mathbf{L}^{\left(q\right)}_{k}\left(r\right)\}$
@@
{adjustwidth}
-0cm
Figure 11: The existence of the multi-order scale-invariance (Part II). (a-e) The degree distributions and the Laplacian eigenvalue spectra of different synthetic interacting systems under the transformation of the SRG are illustrated after being averaged across all replicas. The high-order path Laplacian operator in Eq. (5) is applied to implement the SRG.
Figure 12: The organizational structure transformed by renormalization flows. The Louvain communities of all data sets are adopted from Fig. 7 directly. (a-c) The SRG is applied to renormalize these systems with $\left(p,q\right)=\left(2,1\right)$. The community label of each renormalized super-unit is determined according to the most frequent community label of the units aggregated into it. The values of average path distance between communities, $\kappa$, are measured before and after renormalization, whose absolute differences (i.e., see color-bars) quantify the deviation degrees from the initial organizational structures caused by renormalization. (d-f) The analysis same as (a-c) is implemented with $\left(p,q\right)=\left(2,1\right)$.
@@
according to Eqs. (25-26) in the main text, we can derive
@@
Appendix H Extended results of the multi-order scale-invariance
@@
In this section, we present the extended results of multi-order scale-invariance verification. Our verification is implemented using the behaviours of the Laplacian eigenvalue spectrum.
@@
We implement our experiments on multiple types of synthetic interacting systems, whose pairwise interactions follow the Barabási-Albert network (BA, $c=1$ or $c=10$), the Watts-Strogatz network (WS, each unit initially has $10$ neighbors and edges are rewired according to a probability of $0.1$), the Erdos-Renyi network (ER, each pair of units share an edge with a probability of $0.01$), the triangular lattice (TL), and the random tree (RT), respectively. Each kind of system consists of $1000$ units and has $1000$ replicas.
@@
As shown in Fig. 10, the interacting systems whose pairwise interactions follow weakly scale-invariant (e.g., BA) and scale-invariant (e.g., TL and RT) structures exhibit multi-order scale invariance because their Laplacian eigenvalue spectra generally follow fixed power-law distributions during renormalization (note that noises in Laplacian eigenvalue spectra are unavoidable due to the numerical instability of eigen-decomposition in sparse matrices). For comparisons, the absence of multi-order scale invariance can be observed in the interacting systems with scale-dependent pairwise interactions (e.g., ER and WS). These results are consistent with our findings in Fig. 5, where the degrees to which the scaling relation holds are higher in scale-invariant and weakly scale-invariant systems (i.e., with smaller deviations) and are lower in other systems. In fact, the departures from the scaling relation of scale-dependent systems arise from the absence of power-law behaviours (e.g., see Fig. 10(b-c)). Apart from the verification of multi-order scale invariance, the results shown in Fig. 10 also support the study of the effects of high-order interactions on power-law behaviours. When the SRG is guided by $2$-order interactions, the Laplacian eigenvalue spectra of the interacting systems whose pairwise interactions follow the ER model seem to exhibit power-law behaviours after one iteration of renormalization (i.e., $k>1$). We speculate that specific scale-invariant structures may emerge in the ER model after some high-order interactions are reduced.
@@
The above findings can also be validated by the observations in Fig. 11, where all settings in the computational experiments are same as those in Fig. 10 except that the high-order path Laplacian operator is applied to design the SRG.
Appendix I On the scaling relation and high-order effects
@@
Appendix J On organizational structure preservation
Here we elaborate on the experiments of verifying whether given organizational structures can be preserved by the SRG.
Before renormalization, the communities of units are automatically detected by the Louvain community detection algorithm [90]. After every iteration of renormalization, units are aggregated into super-units according to the SRG. Here we do not define latent community structures by applying the algorithm [90] on the renormalized system again. Instead, we determine latent community structures according to unit aggregation during renormalization, where the community label of each super-unit is defined as the most frequent community label of the units aggregated into it. For instance, if six units are aggregated into a super-unit during renormalization and four of them belong to the first community, then the super-unit is defined as a member of the first community.
@@
Appendix K Programmatic implementation of the SRG
@@
1## Dependency libraries used for the RRG:
@@
In application, we have a system, G, to process. For the SRG, we need to ensure that G is a graph object in the $\mathbf{networkx}$ library. The SRG renormalizes the system on the $p$-order based on $q$-order interactions ($p\leq q$), where $p$ and $q$ can be defined by specifying p and q. Meanwhile, to select the type of the Laplacian representation (i.e., the Multi-order Laplacian operator or the high-order path Laplacian in Sec.II), we need to specify L_Type in the program. Given these settings, the SRG runs on the system for IterNum iterations to generate a renormalization flow.
@@
3 L=HighOrderLaplician(A, L_Type, Order=q)
4 L0=HighOrderLaplician(A, L_Type, Order=p)
5 L_List,L0_List,C_List,Tracked_Alignment=SRG_Function(L,L0,q,IterNum)
6 return L_List,L0_List,C_List,Tracked_Alignment
The main function of the SRG generates four outputs after computation. The first two outputs, L_List and L0_List, are the lists of operator $\mathbf{L}^{\left(q\right)}_{k}$ and operator $\mathbf{L}^{\left(p\right)}_{k}$, respectively. For instance, the first element of L_List is $\mathbf{L}^{\left(q\right)}_{1}$, the second element is $\mathbf{L}^{\left(q\right)}_{2}$, and so on. The number of elements in L_List and L0_List is determined by IterNum.
The third output, C_List, is the list of specific heat $X_{1}^{\left(q\right)}$ vector calculated by the initial $q$-order Laplacian, L_List[0]. The number of specific heat vector in C_List is determined by the number of connected clusters in L_List[0]. For instance, in the ergodicity case, C_List contains only one element, which is a vector of specific heat derived on the only one connected cluster.
The last output of the main function is Tracked_Alignment, which indicates the indexes of the initial units aggregated into each macro-unit in every connected cluster after the $k$-th iteration of the SRG. Below, we present a simple instance where system G contains only six units and satisfies the ergodicity.

### v7

@@ Appendix E The behaviour of the specific heat
As mentioned in our main text and explained in Appendix D, the specific heat is used to define the cutoff of the spectrum of operator $\mathbf{L}^{\left(q\right)}$, which helps determine relevant scales in the system. Specifically, the $q$-order specific heat in Eq. (25), derived using the first derivative of the $q$-order spectral entropy, is used to define the optimal time scale, $\widehat{\tau}^{\left(q\right)}$, of the SRG in Eq. (26). Then, the inverse of time scale $\widehat{\tau}^{\left(q\right)}$ is utilized to help define the number of characteristic modes for describing the system near criticality (see step (4) in the SRG).
@@
Appendix F Bias correction in time scale calculation
To understand the bias of time scale calculation, we need to recall the derivation process of $\widetilde{\tau}^{\left(q\right)}$, a time scale under the infinite deceleration condition
$\frac{\mathsf{d}}{\mathsf{d}\tau^{\left(q\right)}}X^{\left(q\right)}_{1}\left(\tau^{\left(q\right)}\right)\Big{|}_{\widetilde{\tau}^{\left(q\right)}}$ $=0,$ (76)
$\frac{\mathsf{d}}{\mathsf{d}\tau^{\left(q\right)}}\left[-\left(\tau^{\left(q\right)}\right)^{2}\frac{\mathsf{d}}{\mathsf{d}\tau^{\left(q\right)}}\big{\langle}\lambda^{\left(q\right)}_{1}\big{\rangle}_{\rho}\right]\bigg{|}_{\widetilde{\tau}^{\left(q\right)}}$ $=0,$ (77)
$\frac{\mathsf{d}}{\mathsf{d}\tau^{\left(q\right)}}\left[-\left(\tau^{\left(q\right)}\right)^{2}\frac{\mathsf{d}}{\mathsf{d}\tau^{\left(q\right)}}\operatorname{tr}\left(\mathbf{L}^{\left(q\right)}_{1}\rho^{\left(q\right)}_{1}\right)\right]\bigg{|}_{\widetilde{\tau}^{\left(q\right)}}$ $=0,$ (78)
where each $\lambda^{\left(q\right)}_{1}$ is an eigenvalue of $\mathbf{L}^{\left(q\right)}_{1}$. Eq. (77) is directly obtained from Eq. (25). As shown by Eqs. (77-78), the selection of $\widetilde{\tau}^{\left(q\right)}$ is fully determined by the Laplacian operator $\mathbf{L}^{\left(q\right)}_{1}$, its eigenvalue spectrum, and its associated density operator $\rho^{\left(q\right)}_{1}$.
Similar to what we have discussed in Sec. IV (i.e., see step (3) in the SRG), the effectiveness of diffusion may not necessarily be reflected by Eqs. (77-78). Specifically, in the density operator $\rho^{\left(q\right)}_{1}$, there exist numerous possibilities for its $\left(i,j\right)$-th element to be smaller than both the $\left(i,i\right)$-th and the $\left(j,j\right)$-th elements. In these cases, the $q$-order interactions between $v_{i}$ and $v_{j}$ are less stronger than the $q$-order self-interactions of these two units. Given a time scale $\tau^{\left(q\right)}$, the accumulations of the $q$-order interactions between $v_{i}$ and $v_{j}$ are less important than the accumulations of the $q$-order self-interactions in determining the behaviours of $v_{i}$ and $v_{j}$. Therefore, we should treat the diffusion process from $v_{i}$ to $v_{j}$ (or vice versa) as less effective. Eqs. (77-78) may neglect these possibilities because they use $\rho^{\left(q\right)}_{1}$ without further modification. This property leads to a bias in time scale calculation.
Figure 10: Bias in time scale calculation. (a) shows $n_{1}^{\left(q\right)}\left(\tau^{\left(q\right)}\right)$ and $\big{|}\chi_{1}^{\left(q\right)}\left(\tau^{\left(q\right)}\right)\big{|}$ as the functions of the time scale $\tau^{\left(q\right)}$ based on the results obtained on the Erdos-Renyi network (ER), the Watts-Strogatz network (WS), the Barabási-Albert network (BA), the random tree (RT), the triangular lattice (TL), and the square lattice (SL), respectively. Each network is initialized with $300$ units. Among these networks, TL and SL can be defined with or without the periodic boundary condition (PBC). The SRG is defined with $\left(p,q\right)=\left(1,1\right)$. (b) shows the same analysis on these networks, where the SRG is defined with $\left(p,q\right)=\left(1,2\right)$. Note that RT is no longer included in our analysis since it contains no $2$-order interaction. In (a-b), ER networks are defined with different connection probabilities. For instance, “ER, 0.1” denotes the case where a ER network is defined with a connection probability of $0.1$. Under each condition, we represent $n_{1}^{\left(q\right)}\left(\tau^{\left(q\right)}\right)$ by a purple line and represent $\big{|}\chi_{1}^{\left(q\right)}\left(\tau^{\left(q\right)}\right)\big{|}$ by a blue line. Meanwhile, we mark $\widetilde{\tau}^{\left(q\right)}$ and $\tau^{\left(q\right)}_{*}$ by a gray bar and a blue bar, respectively. Therefore, we have $\tau^{\left(q\right)}_{*}>\widetilde{\tau}^{\left(q\right)}$ if the blue bar occurs on the right side of the gray bar, which indicates a bias in time scale calculation.
To explore when this bias occurs, we use $n_{1}^{\left(q\right)}\left(\tau^{\left(q\right)}\right)$, the number of Kadanoff blocks in $\mathbf{M}^{\left(q\right)}_{1}\left(\tau^{\left(q\right)}\right)$, as an observable. Here $\mathbf{M}^{\left(q\right)}_{1}\left(\tau^{\left(q\right)}\right)$ is a reference network derived in a way similar to the network $\mathbf{H}^{\left(q\right)}_{k}$ shown in the step (2) of the SRG in Sec. IV. Specifically, we initialize $\mathbf{M}^{\left(q\right)}_{1}\left(\tau^{\left(q\right)}\right)$ as a null network and calculate
$\rho^{\left(q\right)}_{1}\left(\tau^{\left(q\right)}\right)=\frac{\sum_{\lambda_{1}^{\left(q\right)}}\exp\left(-\tau^{\left(q\right)}\lambda_{1}^{\left(q\right)}\right)\big{|}\lambda_{1}^{\left(q\right)}\big{\rangle}\big{\langle}\lambda_{1}^{\left(q\right)}\big{|}}{\sum_{\lambda_{1}^{\left(q\right)}}\exp\left(-\tau^{\left(q\right)}\lambda_{1}^{\left(q\right)}\right)}.$ (79)
We go through every element in $\rho^{\left(q\right)}_{1}\left(\tau^{\left(q\right)}\right)$. If the $\left(i,j\right)$-th element is greater than the $\left(i,i\right)$-th or the $\left(j,j\right)$-th element, we add an edge between units $v_{i}$ and $v_{j}$ in $\mathbf{M}^{\left(q\right)}_{1}\left(\tau^{\left(q\right)}\right)$. We count the number of clusters in $\mathbf{M}^{\left(q\right)}_{1}\left(\tau^{\left(q\right)}\right)$ as $n_{1}^{\left(q\right)}\left(\tau^{\left(q\right)}\right)$. To reflect how $n_{1}^{\left(q\right)}\left(\tau^{\left(q\right)}\right)$ changes with $\tau^{\left(q\right)}$, we measure the absolute value of a susceptibility-like quantity
$\Big{|}\chi_{1}^{\left(q\right)}\left(\tau^{\left(q\right)}\right)\Big{|}=\bigg{|}\frac{\partial}{\partial\tau}n_{1}^{\left(q\right)}\left(\tau^{\left(q\right)}\right)\bigg{|}_{\tau^{\left(q\right)}}\bigg{|}.$ (80)
As $\tau^{\left(q\right)}$ increases, there is a higher possibility for $n_{1}^{\left(q\right)}\left(\tau^{\left(q\right)}\right)$ to decrease (note that $n_{1}^{\left(q\right)}\left(\tau^{\left(q\right)}\right)$ is not strictly decreasing) since we progressively add more edges into $\mathbf{M}^{\left(q\right)}_{1}\left(\tau^{\left(q\right)}\right)$.
In Fig. 10, we use the results obtained on several representative random network models to show $n_{1}^{\left(q\right)}\left(\tau^{\left(q\right)}\right)$ and $\big{|}\chi_{1}^{\left(q\right)}\left(\tau^{\left(q\right)}\right)\big{|}$ as the functions of the time scale $\tau^{\left(q\right)}$. Meanwhile, we show the time scale $\widetilde{\tau}^{\left(q\right)}$ selected based on Eqs. (77-78) as a reference. In our analysis, we focus on analyzing the relation between $\widetilde{\tau}^{\left(q\right)}$ and $\tau^{\left(q\right)}_{*}$, where $\tau^{\left(q\right)}_{*}$ denotes the time scale that maximizes $\big{|}\chi_{1}^{\left(q\right)}\left(\tau^{\left(q\right)}\right)\big{|}$. As shown in Fig. 10, a time scale $\tau^{\left(q\right)}$ below $\tau^{\left(q\right)}_{*}$ generally corresponds to a null network $\mathbf{M}^{\left(q\right)}_{1}\left(\tau^{\left(q\right)}\right)$ where there exists no effective diffusion. Once the time scale $\tau^{\left(q\right)}$ exceeds $\tau^{\left(q\right)}_{*}$, the value of $n_{1}^{\left(q\right)}\left(\tau^{\left(q\right)}\right)$ exhibits significant reductions on all random network models. Therefore, the time scale $\tau^{\left(q\right)}_{*}$ can serve as an indicator of the emergence of effective diffusion. When we compare between $\widetilde{\tau}^{\left(q\right)}$ and $\tau^{\left(q\right)}_{*}$, we can see that $\widetilde{\tau}^{\left(q\right)}$ is larger than $\tau^{\left(q\right)}_{*}$ in most random network models except for the Erdos-Renyi network (ER) under some conditions. This phenomenon occurs on both pairwise and $2$-order interactions. Consequently, in most random network models, the time scale $\widetilde{\tau}^{\left(q\right)}$ selected according to the infinite deceleration condition shown in Eqs. (77-78) is sufficient to create effective diffusion. As for the ER network, we define it with different connection probability (i.e., two units share an edge with a given probability). When this probability enlarges, the ER network becomes increasingly dense. It can be seen that $\widetilde{\tau}^{\left(q\right)}$ gradually becomes increasingly smaller than $\tau^{\left(q\right)}_{*}$ as the system becomes more dense. A value of $\widetilde{\tau}^{\left(q\right)}$ that is smaller than $\tau^{\left(q\right)}_{*}$ inevitably leads to a bias because there is no effective diffusion included while measuring the specific heat in Eq. (76). To avoid this bias, we suggest a practical way to re-scale $\widetilde{\tau}^{\left(q\right)}$ when it is smaller than $\tau^{\left(q\right)}_{*}$. Specifically, we define a simple coefficient
$\varepsilon=\frac{\tau^{\left(q\right)}_{*}}{\min\{\widetilde{\tau}^{\left(q\right)},\tau^{\left(q\right)}_{*}\}}$ (81)
and always re-scale $\widetilde{\tau}^{\left(q\right)}$ as $\varepsilon\widetilde{\tau}^{\left(q\right)}$ (i.e., if $\widetilde{\tau}^{\left(q\right)}$ is smaller than $\tau^{\left(q\right)}_{*}$, we re-scale $\widetilde{\tau}^{\left(q\right)}$ to $\tau^{\left(q\right)}_{*}$ directly. If $\widetilde{\tau}^{\left(q\right)}$ is no less than $\tau^{\left(q\right)}_{*}$, there is no further adjustment). This re-scaling procedure leads to Eq. (26) in Sec. IV.
Apart from the bias in time scale calculation, Fig. 10 also conveys an important message about the effects of the periodic boundary condition. It can be seen that $n_{1}^{\left(q\right)}\left(\tau^{\left(q\right)}\right)$ generally maintains constant under the periodic boundary condition (there are only small perturbations caused by numerical errors). This is because all units and interactions become identical and the effective diffusion no longer exists. These results provide an explanation for the exactly invariant property of the lattice systems defined with the periodic boundary condition during renormalization. Note that the selection of $\tau^{\left(q\right)}_{*}$ under the periodic boundary condition, as shown by Fig. 10, is mainly determined by the occasional and small numerical errors while measuring $n_{1}^{\left(q\right)}\left(\tau^{\left(q\right)}\right)$.
Appendix G Comparison of the SRG across different settings of $\left(p,q\right)$
In this section, we compare the SRG across different settings of $\left(p,q\right)$ to validate the meaning of distinguishing between $p$ and $q$ during renormalization. In Fig. 11, we use the SRG to process a system whose pairwise interactions follow the Watts-Strogatz network (WS), a random network model exhibiting small-world properties [111].
In our experiment, we renormalize the system on the $p$-order under the guidance of $q$-order interactions. We first consider a case with $p=q$ (i.e., we renormalize the system on an order according to its properties on this order). When $p=q=1$, our analysis becomes similar with other conventional renormalization group frameworks that focus on pairwise interactions. When $p=q>1$, the SRG is implemented on high-order interactions but our analysis involves no cross-order information since we do not distinguish between $p$ and $q$ during renormalization. In these cases, the SRG is only equipped with the information of a single order.
Figure 11: The simplex renormalization group (SRG) with different $\left(p,q\right)$. We generate a system with $200$ units whose pairwise interactions follow a Watts-Strogatz network (each unit initially has $6$ neighbors and edges are rewired according to a probability of $0.05$) [111]. Then, the SRG is applied to renormalize the system on the $p$-order according to the properties of $q$-order interactions. (a-d) show the renormalizations flows in both moment and real spaces, which are generated by the SRG defined with $\left(p,q\right)=\left(1,1\right)$, $\left(p,q\right)=\left(2,2\right)$, $\left(p,q\right)=\left(1,2\right)$, and $\left(p,q\right)=\left(2,3\right)$, respectively.
For comparison, we distinguish between $p$ and $q$ and ensure that $p\neq q$. Thus, we create a possibility for analyzing the system based on cross-order information. Specifically, we consider $\left(p,q\right)=\left(1,2\right)$ and $\left(p,q\right)=\left(2,3\right)$ as two instances. These two cases have not been studied by existing renormalization group frameworks, allowing us to see the nontrivial effects of the interactions on a higher order on the interactions on a lower order.
Our results are presented in Fig. 11, where we illustrate the renormalization flows in both real (i.e., $\mathbf{G}^{\left(p\right)}_{k}$) and moment (i.e., $\mathbf{L}^{\left(p\right)}_{k}$) spaces. By comparing between Figs. 11(a-b) and Figs. 11(c-d), we can directly observe systematic differences between the case with $p=q$ and the case with $p\neq q$. Distinguishing between $p$ and $q$ during renormalization is meaningful and effective, enabling us to explore more general cases uncovered by classic frameworks. Moreover, although high-order interactions are usually distributed in a sparser manner, renormalizing them frequently leads to more drastic system reductions. This is because high-order interactions involve with more units than low-order ones.
@@ Appendix G Comparison of the SRG across different settings of $\left(p,q\right)$
Appendix H Scaling relation of the Laplacian eigenvalue spectrum in the ergodic case
@@
{adjustwidth}
-0cm
Figure 12: The existence of the multi-order scale-invariance (Part I). (a-f) The degree distributions and the Laplacian eigenvalue spectra of different synthetic interacting systems under the transformation of the SRG are illustrated after being averaged across all replicas. The multi-order Laplacian operator in Eq. (1) is applied to implement the SRG. Note that the original Laplacian eigenvalue spectra of all triangular lattices (TL) are calculated using the analytic formula developed by Ref. [113].
{adjustwidth}
-0cm
Figure 13: The existence of the multi-order scale-invariance (Part II). (a-f) The degree distributions and the Laplacian eigenvalue spectra of different synthetic interacting systems under the transformation of the SRG are illustrated after being averaged across all replicas. The high-order path Laplacian operator in Eq. (5) is applied to implement the SRG. Note that the original Laplacian eigenvalue spectra of all triangular lattices (TL) are calculated using the analytic formula developed by Ref. [113].
Appendix I Scaling relation of the Laplacian eigenvalue spectrum in the non-ergodic case
@@
where $\Lambda^{\left(q\right)}_{k}$ and $\Lambda^{\left(q\right)}_{k}\left(i\right)$ denote the eigenvalue sets of $\mathbf{L}^{\left(q\right)}_{k}$ and $\mathbf{L}^{\left(q\right)}_{k}\left(i\right)$, respectively. According to Eqs. (89-90), the characteristic polynomial roots of $\mathbf{L}^{\left(q\right)}_{k}$ can be constructed using the combination of the characteristic polynomial roots of $\{\mathbf{L}^{\left(q\right)}_{k}\left(1\right),\ldots,\mathbf{L}^{\left(q\right)}_{k}\left(r\right)\}$
@@
according to Eqs. (27-28) in the main text, we can derive
@@
Appendix J Extended results of the multi-order scale-invariance
@@
In this section, we present the extended results of multi-order scale-invariance verification. Our verification is implemented using the behaviours of Laplacian eigenvalue spectra and degree distributions.
@@
We implement our experiments on multiple types of synthetic interacting systems, whose pairwise interactions follow the Barabási-Albert network (BA, $c=1$ or $c=2$), the Watts-Strogatz network (WS, each unit initially has $10$ neighbors and edges are rewired according to a probability of $0.1$), the Erdos-Renyi network (ER, each pair of units share an edge with a probability of $0.01$), the triangular lattice (TL), and the random tree (RT), respectively. Each kind of system consists of $1000$ units and has $100$ replicas. We use the SRG defined with different settings of $\left(p,q\right)$ to process these systems on different orders. Based on the obtained renormalization flows, we calculate eigenvalue spectra and degree distributions using $\mathbf{L}^{\left(p\right)}_{k}$ and $\mathbf{G}^{\left(p\right)}_{k}$. Please note that the degree distribution of $\mathbf{G}^{\left(p\right)}_{k}$ gradually becomes weighted during renormalization.
@@
In Fig. 12, we show the results generated by the SRG defined with the multi-order Laplacian operator (MOL). As shown in our results, the interacting systems whose pairwise interactions follow weakly scale-invariant (e.g., BA) and scale-invariant (e.g., TL and RT) structures exhibit multi-order scale invariance because their Laplacian eigenvalue spectra generally follow fixed power-law distributions during renormalization (note that noises in Laplacian eigenvalue spectra are unavoidable due to the numerical instability of eigen-decomposition in sparse matrices). For comparisons, the absence of multi-order scale invariance can be observed in the interacting systems with scale-dependent pairwise interactions (e.g., ER and WS). These results are consistent with our findings in Fig. 6, where the degrees to which the scaling relation holds are higher in scale-invariant and weakly scale-invariant systems (i.e., with smaller deviations) and are lower in other systems. In fact, the departures from the scaling relation of scale-dependent systems arise from the absence of power-law behaviours (e.g., see Fig. 12(b-c)).
Apart from the verification of multi-order scale invariance, the results shown in Fig. 12 also support the study of the effects of high-order interactions on power-law behaviours. When the SRG is guided by $2$-order interactions, the Laplacian eigenvalue spectra of the interacting systems whose pairwise interactions follow the ER model seem to exhibit power-law behaviours after one iteration of renormalization (i.e., $k>1$). We speculate that specific scale-invariant structures may emerge in the ER model after some high-order interactions are reduced.
@@
The above findings can also be validated by the observations in Fig. 13, where all settings in the computational experiments are same as those in Fig. 12 except that the high-order path Laplacian operator (HOPL) is applied to design the SRG.
Figure 14: The organizational structure transformed by renormalization flows. The Louvain communities of all data sets are adopted from Fig. 8 directly. (a-c) The SRG is applied to renormalize these systems with $\left(p,q\right)=\left(2,1\right)$. The community label of each renormalized macro-unit is determined according to the most frequent community label of the units aggregated into it. The values of average path distance between communities, $\kappa$, are measured before and after renormalization, whose absolute differences (i.e., see color-bars) quantify the deviation degrees from the initial organizational structures caused by renormalization. (d-f) The analysis same as (a-c) is implemented with $\left(p,q\right)=\left(2,1\right)$.
Appendix K On the scaling relation and high-order effects
@@
Appendix L On organizational structure preservation
Here we elaborate on the experiments of verifying whether given organizational structures can be preserved by the SRG framework.
Before renormalization, the communities of units are automatically detected by the Louvain community detection algorithm [90]. After every iteration of renormalization, units are aggregated into macro-units according to the SRG. Here we do not define latent community structures by applying the algorithm [90] on the renormalized system again. Instead, we determine latent community structures according to unit aggregation during renormalization, where the community label of each macro-unit is defined as the most frequent community label of the units aggregated into it. For instance, if six units are aggregated into a macro-unit during renormalization and four of them belong to the first community, then the macro-unit is defined as a member of the first community.
@@
Appendix M Programmatic implementation of the SRG
@@
1## Dependency libraries used for the SRG:
@@
In application, we have a system, G, to process. For the SRG, we need to ensure that G is a graph object in the $\mathbf{networkx}$ library. The SRG renormalizes the system on the $p$-order based on the $q$-order interactions ($p\leq q$), where $p$ and $q$ can be defined by specifying p and q. Meanwhile, to select the type of the Laplacian representation (i.e., the Multi-order Laplacian operator or the high-order path Laplacian in Sec.II), we need to specify L_Type in the program. Given these settings, the SRG runs on the system for IterNum iterations to generate a renormalization flow.
@@
3 Lq=HighOrderLaplician(A, L_Type, Order=q)
4 Lp=HighOrderLaplician(A, L_Type, Order=p)
5 Lq_List,Lp_List,Gq_List,Gp_List,C_List,Tracked_Alignment=SRG_Function(Lq,Lp,q,IterNum)
6 return Lq_List,Lp_List,Gq_List,Gp_List,C_List,Tracked_Alignment
The main function of the SRG generates six outputs after computation. The first two outputs, Lq_List and Lp_List, are the lists of operator $\mathbf{L}^{\left(q\right)}_{k}$ and operator $\mathbf{L}^{\left(p\right)}_{k}$, respectively. For instance, the first element of Lq_List is $\mathbf{L}^{\left(q\right)}_{1}$, the second element is $\mathbf{L}^{\left(q\right)}_{2}$, and so on. The number of elements in Lq_List and Lp_List is determined by IterNum.
The third and forth outputs, Gq_List and Gp_List, are the lists of the associated network sketches $\mathbf{G}^{\left(q\right)}_{k}$ and $\mathbf{G}^{\left(p\right)}_{k}$ of the above Laplacian operators, respectively. For example, the first element of Gq_List is $\mathbf{G}^{\left(q\right)}_{1}$, the second element is $\mathbf{G}^{\left(q\right)}_{2}$, and so on. The number of elements in Gq_List and Gp_List equals to IterNum.
The fifth output, C_List, is the list of specific heat $X_{1}^{\left(q\right)}$ vector calculated by the initial $q$-order Laplacian, Lq_List[0]. The number of specific heat vector in C_List is determined by the number of connected clusters in Lq_List[0]. For instance, in the ergodicity case, C_List contains only one element, which is a vector of specific heat derived on the only one connected cluster.
The last output of the main function is Tracked_Alignment, which indicates the indexes of the initial units aggregated into each macro-unit in every connected cluster after the $k$-th iteration of the SRG. Below, we present a simple instance where the system G contains only six units and satisfies the ergodicity.